OpenStudy (anonymous):

a vintage car was bought for 30000 dollars and was sold 10 years later for 50,000 dollars. find the effective yield?

OpenStudy (anonymous):

do you know the formula to find the interest rate or the yield rate?

OpenStudy (anonymous):

interest ya simple or compound

OpenStudy (anonymous):

simple

OpenStudy (anonymous):

a=p(1+r)^t

OpenStudy (anonymous):

okay great! so what is a , p and t in your problem?

OpenStudy (anonymous):

idk what effective yield is and thats what im finding

OpenStudy (anonymous):

effective yield is nothing but the interest rate

OpenStudy (anonymous):

oh ok all i know is the question i asked for the prolem

OpenStudy (anonymous):

effective yield is the effective annual interest rate.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Effective_interest_rate

OpenStudy (anonymous):

Effective Rate (Effective Yield) The effective rate is the actual rate that you earn on an investment or pay on a loan after the effects of compounding frequency are considered. To make a fair comparison between two interest rates when different compounding periods are used, you should first convert both nominal (or stated) rates to their equivalent effective rates so the effects of compounding can be clearly seen. The effective rate of an investment will always be higher than the nominal or stated interest rate when interest is compounded more than once per year. As the number of compounding periods increases, the difference between the nominal and effective rates will also increase. To convert a nominal rate to an equivalent effective rate: Effective Rate = (1 + (i / n))n - 1 Where: i = Nominal or stated interest rate n = Number of compounding periods per year Example: What effective rate will a stated annual rate of 6% yield when compounded semiannually? Effective Rate = ( 1 + .06 / 2 )2 - 1 = .0609

OpenStudy (anonymous):

But i think in your problem, effective rate is just the simple interest, because there is no mention of it being compounded. So just plug in the values in a=p(1+r)^t and get r.